CH​ (g) + 2O​ (g) → CO​ (g) + 2H​ O(g), if 15.00 L of CH​ is burned with 4​ 2​ 2​ 2​ 4​
excess O​ at 23°C and 1.05 atm, what mass of H​ O is produced
1
Expert's answer
2020-04-23T13:07:42-0400

P = 1.05 atm;

V = 15.00 L;

T = 23oC = (23oC + 273.15) = 296.15 K;

R = 0.08206 L atm K-1 mol-1.


Solution:

The balanced chemical equation is:

CH​4(g) + 2O​2(g) → CO​2(g) + 2H​2O(g)

Let's use the ideal gas equation to find the moles of CH4.

The ideal gas equation can be expressed as: PV = nRT.

n = PV / RT

Then,

n(CH4) = (1.05 atm × 15.00 L) / (0.08206 L atm K-1 mol-1 × 296.15 K) = 0.6481 mol

n(CH4) = 0.6481 mol.


According to the chemical equation: n(CH4) = n(H2O)/2

n(H2O) = 2 × n(CH4) = 2 × 0.6481 mol = 1.2962 mol.


Moles of H2O = Mass of H2O / Molar mass of H2O

Molar mass of H2O = M(H2O) = 18.01528 g/mol.

Finally, we can find the number of grams of H2O:

n(H2O) = m(H2O) / M(H2O);

m(H2O) = n(H2O) × M(H2O);

m(H2O) = (1.2962 mol × 18.01528 g/mol) = 23.3514 g = 23.35 g

Mass of H2O = m(H2O) = 23.35 g.


Answer: 23.35 grams of H2O is produced.